The physical fitness of an athlete is often measured by how much oxygen the athlete takes in (which is recorded in milliliters per kilogram, ml/kg). The mean maximum oxygen uptake for elite athletes has been found to be 66 with a standard deviation of 7.4. Assume that the distribution is approximately normal.

Find the probability that an elite athlete has a maximum oxygen uptake of at least 77.8 ml/kg.

Find the probability that an elite athlete has a maximum oxygen uptake of at most 49.7 ml/kg
Answer #1

Solution :

Given that ,

mean = = 66

standard deviation = = 7.4

1) P(x 77.8) = 1 - P(x   77.8)

= 1 - P[(x - ) / ( 77.8 - 66 ) / 7.4 ]

= 1 -  P(z 1.59)   

  Using z table,

= 1 - 0.9441

= 0.0559

2) P(x 49.7 )

= P[(x - ) / ( 49.7 - 66 ) / 7.4 ]

= P(z -2.20)

Using z table,

= 0.0139